40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 16 03:16:50 2009, rana, rob, Summary, LSC, funny timing setup on the LSC 1X3_1.JPG
    Reply  Fri Oct 16 13:25:18 2009, Koji, Summary, LSC, funny timing setup on the LSC 
    Reply  Mon Oct 19 14:48:15 2009, rana, rob, Summary, Electronics, piezo jena measuring box PA160153.JPGPA160151.JPG
Message ID: 2101     Entry time: Fri Oct 16 03:16:50 2009     Reply to this: 2104   2118
Author: rana, rob 
Type: Summary 
Category: LSC 
Subject: funny timing setup on the LSC 

While measuring the Piezo Jena noise tonight we noticed that the LSC timing is setup strangely.

Instead of using the Fiber Optic Sander Liu Timing board, we are just using a long 4-pin LEMO cable which comes from somewhere in the cable tray. This is apparent in the rack pictures (1X3) that Kiwamu has recently posted in the Electronics Wiki. I think all of our front ends are supposed to use the fiber card for this. I will ask Jay and Alex what the deal is here - seems like to me that this can be a cause for timing noise on the LSC.

We should be able to diagnose timing noise between the OMC and the LSC by putting in a signal in the OMC and looking at the signal on the LSC side. Should be a matlab script that we can run whenever we are suspicious of this. This is an excellent task for a new visiting grad student to help learn how to debug the digital control system.

Attachment 1: 1X3_1.JPG  3.620 MB  | Hide | Hide all
1X3_1.JPG
ELOG V3.1.3-